# CounterCurate: Enhancing Physical and Semantic Visio-Linguistic Compositional Reasoning via Counterfactual Examples

### 20 Feb 2024 v1

@@ 3.5 Fine-tuning Methodology
As demonstrated in Figure 3, CounterCurate can be used to fine-tune both contrastive Radford et al. (2021) and generative Liu et al. (2023b) multimodal models, which significantly improve their performance as shown in experiments. To fine-tune a contrastive model like CLIP, we sample $N$ positive image-text pairs and the corresponding negative image-text pairs to form a batch, as shown in Figure 3 (a). We refer to this method as Grouping. In this way, our model is forced to distinguish the positive pairs from their corresponding negatives, which is demonstrated to be helpful in experiments. We also conduct a detailed analysis of the effectiveness of this method in Section 4.5. The training loss is the same as the original, where cross-entropy loss is used to maximize the diagonal elements and minimize all other entries in the similarity matrix.
@@ 4 Experiments
We utilize the proposed three datasets, Flickr30k-Positions, Flickr30k-Counting, and Flickr30k-Attributes, created via CounterCurate, to fine-tune both contrastive and generative models. Specifically, we select two common multimodal models, ViT-B/32 from Ilharco et al. (2021) OpenCLIP pre-trained on LAION-2B Schuhmann et al. (2022), and LLaVA-1.5 Liu et al. (2023a) as the base model.
@@ 4.3 Evaluating Object Counting
Here we show the counting capability of our models fine-tuned on the Flickr30k-Counting dataset. We select a different dataset, PointQA-LookTwice dataset Mani et al. (2022), as the evaluation benchmark. Specifically, PointQA-LookTwice is designed to ask models the number $n_{j}$ of occurrences of an object $j$ in an image $I$. We reformat this dataset such that for every $j\in I$ with object name $c_{j}$, we generate a positive caption $C_{j}=$ “There are $n_{j}$ $c_{j}$ and a negative caption $C_{j}^{\prime}=$ “There are $n_{j}+1$ $c_{j}$. For example, given the positive caption “there are 3 dogs”, the negative caption is “there are 4 dogs”. We use a similar evaluation protocol as Section 4.2. We mark that a model made a correct prediction if (i) CLIP shows $s_{(C_{j},I)}>s_{(C_{j}^{\prime},I)}$ and (ii) LLaVA-1.5 correctly chooses the option for $C_{j}$. The results are shown in Table 2.
@@
As CLIP performs slightly better than random guessing, it is surprising that LLaVA-1.5 performs worse than random. Nevertheless, fine-tuning with Flickr30k-Counting improves both models’ counting capability. This shows the effectiveness of using GLIGEN-generated negative images in CounterCurate to tackle the problem of counting. Further ablations on removing the negative image-caption pairs or without using grouping can be found in Appendix F.
@@ 4.4 Evaluating Semantic Compositional Reasoning
Similar to the setup in Sec 4.3, we fine-tune both CLIP and LLaVA-1.5 under our curated Flickr30k-Attributes. We use another common benchmark SugarCrepe Hsieh et al. (2023) to evaluate the model’s semantic compositional capabilities. The benchmark has three major categories as depicted in Table 3. The evaluation protocol is also the same as Sec 4.3. We choose NegCLIP Yuksekgonul et al. (2023) and GPT-4V OpenAI (2023c) as the representative strong baselines for contrastive and generative models respectively.
We observe significant improvements for both CLIP and LLaVA-1.5, both on average and categorically. The main results are shown in Table 3. For detailed scores for each subcategory in SugarCrepe, please see Appendix G.
For example, CounterCurate fine-tuned CLIP model surpasses NegCLIP on average as well as in two main categories. Note that the source of fine-tuning data, Flickr30k-Entities, contains much fewer image-caption pairs than COCO-Captions (around 100k image-caption pairs) where NegCLIP is trained. Furthermore, when prompting GPT-4V to generate the negative captions, we intentionally prompted GPT-4V to produce None when there is no feasible condition to conduct “swapping". Such two factors lead to the fact that our data curation pipeline results much fewer negative samples for the “swap” category, which is around 2.5k. We argue that using our pipeline on other datasets that can generate more of the “swap” attribute shall lead to larger improvements in performance.
LLaVA-1.5 also performed better in all three categories after fine-tuning. It is also surprising that our fine-tuned model outperforms the SOTA LMM GPT-4V both on average and in two categories, the most significant boost over the “add” category. We observe improvements on other datasets Diwan et al. (2022) as well, attached in Appendix E.
@@ 4.5 In-Depth Analysis
As shown in Table 4, even though using either the negative caption or the negative image can improve performance, the scores are significantly worse than when using both elements to fine-tune. This demonstrates that CounterCurate, which incorporates both negative captions and negative images, is necessary to achieve desirable improvements. It also shows that using grouping allows further improvements compared to not using this strategy.
@@ Acknowledgement
This work was supported in part by NSF CAREER IIS2150012, and Institute of Information & communications Technology Planning & Evaluation(IITP) grants funded by the Korea government(MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No. RS2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training).
@@
We acknowledge that DALLE-3, one of the generative models we use, is a closed-source model, which makes it difficult to systematically analyze its behavior.

### 12 Mar 2024 v2

@@ 3.5 Fine-tuning Methodology
As demonstrated in Figure 3, CounterCurate can be used to fine-tune both contrastive Radford et al. (2021) and generative Liu et al. (2023b) multimodal models, which significantly improve their performance. To fine-tune a contrastive model like CLIP, we sample $N$ positive image-text pairs and the corresponding negative image-text pairs to form a batch $(C,I,C^{\prime},I^{\prime})$ as shown in Figure 3 (a). We refer to this method as Grouping, where our model is forced to distinguish the positive pairs from their corresponding negatives, which is demonstrated to be helpful in experiments. We also conduct a detailed analysis of the effectiveness of this method in Section 4.5. The training loss is the same as the original, where cross-entropy loss is used to maximize the diagonal elements and minimize all other entries in the similarity matrix.
@@ 4 Experiments
Figure 5: Qualitative examples of models’ compositional reasoning capabilities before/after being finetuned via our approach CounterCurate. Wrong answers are marked in red. Our approach enhances both CLIP and LLaVA’s reasoning capabilities.
We utilize the proposed three datasets, Flickr30k-Positions, Flickr30k-Counting, and Flickr30k-Attributes, created via CounterCurate, to fine-tune both contrastive and generative models. Specifically, we select two common multimodal models, ViT-B/32 from Ilharco et al. (2021) OpenCLIP pre-trained on LAION-2B Schuhmann et al. (2022) and LLaVA-1.5 Liu et al. (2023a) as base models. An overview of our results is attached in Figure 5.
@@ 4.3 Evaluating Object Counting
Here we show the counting capability of our models fine-tuned on Flickr30k-Counting. We select a different dataset, PointQA-LookTwice Mani et al. (2022), as the evaluation benchmark. Specifically, PointQA-LookTwice was designed to ask models the number $n_{J}$ of occurrences of an object category $J$ in an image $I$. We reformat this dataset such that for every $J$, we generate a positive caption $C_{J}=$ “There are $n_{J}$ $J$’s” and a negative caption $C_{J}^{\prime}=$ “There are $n_{J}+1$ $J$’s”. For example, given $C_{J}$ “there are 3 dogs”, $C_{J}^{\prime}$ is “there are 4 dogs”. Similar to what we did in Section 4.2, we mark that a model made a correct prediction if (i) CLIP shows $s_{(C_{J},I)}>s_{(C_{J}^{\prime},I)}$ and (ii) LLaVA-1.5 correctly chooses the option for $C_{J}$. The results are shown in Table 2.
@@
As CLIP performs slightly better than random guessing, it is surprising that LLaVA-1.5 performs worse than random. Nevertheless, fine-tuning with Flickr30k-Counting improves both models’ counting capability. This shows the effectiveness of using GLIGEN-generated negative images in CounterCurate to tackle the problem of counting. We conduct further ablation studies in Appendix F.
@@ 4.4 Evaluating Semantic Compositional Reasoning
Similar to the setup in Sec 4.3, we fine-tune both CLIP and LLaVA-1.5 under our curated Flickr30k-Attributes. We use another common benchmark SugarCrepe Hsieh et al. (2023) to evaluate the model’s semantic compositional reasoning capabilities. SugarCrepe has three major categories depicted in Table 3. The evaluation protocol is also the same as Sec 4.3. We choose NegCLIP Yuksekgonul et al. (2023) and GPT-4V OpenAI (2023c) as the representative strong baselines for contrastive and generative models respectively.
We observe significant improvements for both CLIP and LLaVA-1.5, both on average and categorically. Summarized results are shown in Table 3, where Appendix G contains more detailed scores.
For example, CounterCurate fine-tuned CLIP surpasses NegCLIP on average as well as in two main categories. Note that the source of fine-tuning data, Flickr30k-Entities, contains much fewer image-caption pairs than COCO-Captions (around 100k image-caption pairs) where NegCLIP is trained. Furthermore, when prompting GPT-4V to generate the negative captions, we intentionally prompted GPT-4V to produce None when there is no feasible condition to conduct “swapping". Such two factors lead to the fact that our data curation pipeline results much fewer negative samples for the “swap” category, which is around 2.5k. We argue that using our pipeline on other datasets that can generate more of the “swap” attribute shall lead to larger improvements in performance.
LLaVA-1.5 also performed better in all three categories after fine-tuning. It is also surprising that our fine-tuned model outperforms the SOTA LMM GPT-4V both on average and in two of the categories, most significantly over the “add” category. We observe improvements on other datasets Diwan et al. (2022) as well in Appendix E.
@@ 4.5 In-Depth Analysis
As shown in Table 4, even though using either the negative caption or the negative image can improve performance, the scores are significantly worse than when using both elements to fine-tune. This demonstrates that CounterCurate, which incorporates both negative captions and negative images, is necessary to achieve desirable improvements. Grouping also delivers further improvements compared to not using this strategy.
@@ Acknowledgement
This work was supported in part by NSF CAREER IIS2150012, and Institute of Information & communications Technology Planning & Evaluation(IITP) grants funded by the Korea government(MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No. RS2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training), and Microsoft Accelerate Foundation Models Research Program.
@@
We acknowledge that DALLE-3, one of the generative models we use, is a closed-source text-to-image generation model, which makes it difficult to systematically analyze its behavior.